# Mixed Quantum–Classical Dynamics Yields Anharmonic Rabi Oscillations

### v1

@@ I Introduction
Mixed quantum–classical (MQC) dynamics offers a powerful framework for the simulation and interpretation of a wide range of phenomena. Throughout the last few decades, MQC dynamics has established itself as a preferred approach for the low-cost simulation of molecular systems, wherein nuclear motion is represented by classical coordinates while a quantum treatment is reserved for the electronic states C. Tully (1998); Nelson et al. (2014); Subotnik et al. (2016); Wang, Akimov, and Prezhdo (2016); Curchod and Martínez (2018); Crespo-Otero and Barbatti (2018). In recent years, it has been similarly employed for the modeling of materials Nie et al. (2014); Smith and Akimov (2019); Jiang et al. (2021); Lively et al. (2024); Krotz and Tempelaar (2024) and condensed-phase correlated phenomena Petrović, Weber, and Freericks (2024).
@@
Although MQC dynamics is commonly found to provide accurate, or at least predictive, results, its realm of applicability remains poorly understood. According to the Bohr correspondence principle Bohr (1920), the classical approximation taken within MQC dynamics is justified so long as the relevant coordinates attain large quantum numbers. However, MQC dynamics has been shown to reach high accuracy even in cases where this criterion is grossly violated Tempelaar and Reichman (2017); Bondarenko and Tempelaar (2023). Such favorable performance may be explained based on the quasiclassical dynamics literature, where it has been shown that classical equations of motion approach quantum dynamics for harmonic potentials and in the short-time limit Heller (1981); Herman and Kluk (1984); Kay (2006). Yet, counterintuitive behaviors may arise once quantum and classical equations of motion are combined within a same framework, as famously exemplified by the Schrödinger cat thought experiment Schrödinger (1935). As an alternative, one may therefore resort to decoherence as a criterion for the applicability of MQC dynamics, meaning that the quantum–classical correlations should remain limited.
@@
While most applications of MQC dynamics involve electron–nuclear coupling, light–matter interactions provide a complementary realm of applicability Sakko, Rossi, and Nieminen (2014); Ruggenthaler et al. (2018); Hoffmann et al. (2019a, b, 2020); Hsieh, Krotz, and Tempelaar (2023a, b); Weight, Li, and Zhang (2023). Notably, the canonical Rabi model, involving a two-level system (TLS) interacting with a single optical mode, was originally proposed within a MQC treatment, invoking a classical approximation for the mode Rabi (1937). While originally investigated under high field intensities and with neglect of feedback of the TLS onto the mode, a fully self-consistent treatment has since been explored A. Cives-esclop and Sánchez-Soto (1999), as well as behaviors in the quantum–classical crossover region Twyeffort Irish and Armour (2022). However, rapid advances in the areas of quantum optics Reiserer and Rempe (2015) and cavity quantum electrodynamics Ebbesen (2016); Ribeiro et al. (2018); Baranov et al. (2018); Hertzog et al. (2019); Hirai, Hutchison, and Uji-i (2020); Dunkelberger et al. (2022); Mandal et al. (2023); Xiang and Xiong (2024) have generated a particular interest in phenomena where field intensities approach the single-photon limit, as governed by the quantum Rabi model. Much remains to be learned about the applicability of MQC dynamics in this limit, and the ability of the quantum Rabi model to expose general guiding principles for MQC modeling.
@@
In this Article, we apply MQC dynamics to the quantum Rabi model, specifically considering a TLS in its excited state interacting with a single resonant optical mode prepared in its ground state. When represented fully quantum-mechanically, this model is known to exhibit Rabi oscillations manifested as harmonic population fluctuations of the TLS Allen and Eberly (2012). As we show analytically within the rotating wave approximation (RWA), upon adopting a classical representation for the optical mode, Rabi oscillatory motion is instead governed by an undamped and unforced Duffing equation Duffing (1918). This equation involves an anharmonic term contributing a “mode softening effect” to the Rabi oscillations. The solutions to this equation are found to follow the quantum results closely upon (approximately) reinforcing zero-point energy when initializing the classical coordinate. These findings inform on the application of MQC dynamics to coherent phenomena involving small quantum numbers.
@@ II Full-quantum representation
Here, $\hat{q}$ and $\hat{p}$ are coordinate-like operators representing the optical mode, |$\ket{\phi_{\mathrm{g}}}$ and |$\ket{\phi_{\mathrm{e}}}$ refer to the ground and excited state of the TLS, $\mu$ is the associated transition dipole, $\lambda$ denotes the coupling strength between the TLS and the mode, and H.c. is short for Hermitian conjugate.
Figure 1: Schematic of the Rabi model. Shown are the ground and excited states of a two-level system (TLS), denoted |$\ket{\phi_{\mathrm{g}}}$ and |$\ket{\phi_{\mathrm{e}}}$, respectively. Also shown are the lowest two quantum levels of an optical mode, superimposed on a classical harmonic oscillator depiction, involving position and momentum coordinates $q$ and $p$, respectively. The classical optical occupancy, $z^{*}z$, is indicated. The coupling between the TLS and the optical mode is denoted as $g$.
The quantum Rabi Hamiltonian, Eq. 1, is commonly formulated by expressing the TLS in terms of Pauli spin matrices. We refrain from doing so, and instead adopt a ground- and excited-state notation, in order to simplify the forthcoming analysis, and to enforce a connection to quantum optics and quantum electrodynamics where the TLS commonly represents the lowest two states of an atom or other quantum emitter. In both research areas, optical phenomena commonly reach the single-quantum limit, while being amenable to the RWA. Under these conditions, Eq. 1 can be simplified by first expanding the position-like operator $\hat{q}$ into the optical ladder operators,
@@
and by introducing the effective coupling constant $g\equiv\mu\lambda\sqrt{\Omega/2}$ (see Fig. 1). Application of the RWA, which relies on $\Omega\gg g$, then yields a single-quantum Jaynes–Cummings (JC) Hamiltonian Jaynes and Cummings (1963); Shore and Knight (1993) of the form
@@
which exposes Rabi oscillations at a frequency of $2g$.
Upon a replacement of the optical mode by a classical oscillator, Rabi oscillations can no longer be conveniently evaluated in the eigenbasis of the quantum Hamiltonian, and are to be evaluated in the physical basis instead. In order to contrast the resulting approach with that taken for the full-quantum representation, it is worth re-evaluating the latter within the physical basis. We reiterate that a formulation of the quantum Rabi Hamiltonian in terms of the TLS ground and excited states, rather than Pauli matrices, provides notational convenience in this evaluation, especially when we restrict ourselves to pure states of the TLS. [^1] We note that mixed states can always be evaluated by taking an average over pure-state trajectories. In this case, we can describe the TLS in terms of the wavefunction
@@
whose evolution is governed by the time-dependent Schrödinger equation, $i\ket{\dot{\Psi}}=\hat{H}\ket{\Psi}$. This yields dynamical equations for the wavefunction coefficients given by
@@
When $\bar{\epsilon}\gg g$ (which is equivalent to $\Omega\gg g$), the evolution of the wavefunction coefficients will be dominated by phase oscillations at a frequency $\bar{\epsilon}$, due to the diagonal contributions to Eq. 9. Notably, these phase oscillations do not contribute to the population dynamics, which is instead dominated by a slower dynamical component. For that reason, it proves convenient to extract this component by dividing out the phase oscillations. Accordingly, we take
@@ III Mixed quantum–classical representation
We now proceed by taking the classical approximation for the optical mode, starting from the quantum Rabi model. Accordingly, we replace the coordinate-like operators by their classical canonical coordinates in Eq. 1, i.e., $\hat{p}\rightarrow p$ and $\hat{q}\rightarrow q$. This substitution yields a MQC Rabi model, in which the the optical mode is represented by a classical harmonic oscillator, as depicted in Fig. 1.222Here, avoid the commonly used “semiclassical” denomination, in order to differentiate such treatment from that of a quasiclassical approach, for which similar terminology is oftentimes used. In the following, it proves convenient to represent the optical mode by means of a complex-valued classical coordinate defined as Krotz, Provazza, and Tempelaar (2021); Miyazaki, Krotz, and Tempelaar (2024)
@@
Accordingly, the MQC Rabi Hamiltonian takes the form
@@
We will now follow an approach similarly to that presented above for the full-quantum representation in the physical basis. As such, this approach differs from that taken previously for the MQC Rabi model based on Pauli matrices A. Cives-esclop and Sánchez-Soto (1999), and even though we sacrifice notational generalizability to mixed states, it simplifies the analysis considerably. First, the dynamical equations for the wavefunction coefficients take the form
@@
As before, it proves convenient to divide out the phase oscillations from the wavefunction coefficients. However, according to Eq. 17, the frequency of these phase oscillations is now dependent on $n$, which itself evolves in time. Accordingly, we take
@@
where we adopted the RWA as we did for the quantum Rabi model. From this, it follows that
@@
So long as $\Omega\gg g$ is satisfied, the total combined energy of the TLS and the optical field is well approximated as $E=N\Omega$, with $N=\left(n+\tilde{c}_{\mathrm{e}}^{*}\tilde{c}_{\mathrm{e}}\right)$ representing the sum of the TLS and the optical occupancies. Since $E$ is a conserved quantity, so is $N$. If we initialize the TLS in its excited state, i.e., $\tilde{c}_{\mathrm{e}}=1$, and denote the initial optical occupancy as $n_{0}$, we thus have that $n+\tilde{c}_{\mathrm{e}}^{2}=n_{0}+1$. Moreover, if $\tilde{c}_{\mathrm{e}}$ initializes as real-valued, it will remain real-valued throughout the dynamics. Consequently, Eq. 29 simplifies to
@@
Eq. 30 takes the form of an undamped and unforced Duffing equation Duffing (1918), with a harmonic term $-(2+n_{0})g^{2}$ and with an anharmonic term $2g^{2}$. The Duffing equation is ubiquitous in many branches of physics and engineering, and has long been under scrutiny Kovacic and Brennan (2011); Cvetićanin (2013). Eq. 30 is generally solved in terms of Jacobi elliptic functions, with specific solutions dependent on the initial conditions taken. We note that insightful simplified forms of these solutions have remained out of reach. What is known, however, is that the solutions are periodic Abramowitz and Stegun (1968). As such, while the Duffing equation governing the population dynamics within the MQC Rabi model is markedly different from the quantum equivalent, Eq. 13, we find persistent oscillations to be retained, with oscillatory behavior dependent on the initial optical occupancy, $n_{0}$. Within this oscillatory motion, the anharmonic term in the Duffing equation exerts a “mode softening” effect.
@@
Incorporation of optical vacuum fluctuations thus require an initialization satisfying $n_{0}>0$. There are various possibilities in which this criterion can be satisfied. For now, however, we chose to remain agnostic to these possibilities, and instead proceed to systematically and numerically assess the effect of classical coordinate initialization on the MQC Rabi oscillations governed by Eq. 30. Notably, through the dependence of this equation on the initial optical occupancy, $n_{0}=z_{0}^{*}z_{0}$, these oscillations depend on the norm of the initial coordinate $z_{0}$, but not on its phase. This is a direct consequence of the condition $\Omega\gg g$, implying that phase oscillations are infinitely rapid compared to the coherent energy exchange between the TLS and the optical mode.
Figure 2: (a) Transient excited-state population of the two-level system (TLS), $P_{\mathrm{e}}=\tilde{c}_{\mathrm{e}}^{2}$, obtained from the solutions of the undamped and unforced Duffing equation, Eq. 30. Shown are results as a function of the initial optical occupancy, $n_{0}$. (b) Fourier transform of $P_{\mathrm{e}}$. Asymptotic frequency of the dominant harmonic given by Eq. 34 (black dash) and quantum Rabi oscillation at $\omega=2g$ (red dash) are indicated.
Shown in Fig. 2 (a) is the transient excited-state population of the TLS, $P_{\mathrm{e}}=\tilde{c}_{\mathrm{e}}^{2}$, obtained from the solutions to Eq. 30, as a function of $n_{0}$. As anticipated, clear periodic oscillations are observed. Evidently, the oscillation frequency is proportional to $n_{0}$, suggesting that the lack of dynamics in the limit $n_{0}\rightarrow 0$ can be understood as an oscillation with infinite periodicity. The frequency trend is further elucidated in Fig. 2 (b), showing the Fourier transform of $P_{\mathrm{e}}$ as a function of frequency, $\omega$, and initial optical occupancy, $n_{0}$. For each $n_{0}$, the Fourier transform was performed over a duration of $gt=200$, after multiplication of $P_{\mathrm{e}}$ by an exponential with an $1/e$ time of $gt=20$ to avoid clipping and after subtracting a 0.5 baseline, upon which the signal was normalized to the maximum peak. A dominant oscillatory component is observable alongside several minor components that vanish as $n_{0}$ increases. This is suggestive of harmonic oscillatory motion being recovered for $n_{0}\gg 0$. This limit turns out to be amenable to an approximate analytical treatment, providing deeper insights into MQC Rabi oscillations.
@@
This asymptote is indicated in Fig. 2, and the dominant component of the calculated MQC Rabi oscillation is seen to indeed approach this frequency with increasing $n_{0}$.
@@
The asymptotic frequency given by Eq. 34 forms an interesting contrast with the known behavior of the Rabi model within a full-quantum treatment, for which we have $\omega=2g\sqrt{n_{0}+1}$ (with $n_{0}$ representing the quantum occupancy). Interestingly, the MQC Rabi model with neglect of feedback of the TLS on the classical optical field instead yields $\omega=2g\sqrt{n_{0}}$ Allen and Eberly (2012). As such, the self-consistent MQC Rabi model presented here emerges as an interesting “intermediate” between those limiting cases.
@@
Another curious aspect of Eq. 34 pertains to the single-quantum limit. Here, Eq. 34 predicts $n_{0}=1/2$ to be the value for which MQC Rabi oscillations reproduce quantum Rabi oscillations at $\omega=2g$. For this occupancy, the classical field assumes an energy corresponding to that of vacuum fluctuations. This suggests that for MQC Rabi oscillations to reproduce the full-quantum result, the optical field needs to be initialized roughly at the zero-point energy. However, as seen in Fig. 2 (b), the dominant MQC Rabi component begins to undershoot Eq. 34 with decreasing occupancy, in order to satisfy $\omega\rightarrow 0$ for $n_{0}\rightarrow 0$. To offset this effect, MQC Rabi oscillations reproduce the quantum result (also indicated in Fig. 2 (b)) most closely at an occupancy exceeding the zero-point value, that is, at approximately $n_{0}=0.59$. This is evident from Fig. 2, where the quantum Rabi oscillation at $\omega=2g$ is indicated.
@@
Figure 3: Excited-state population of the TLS, $P_{\mathrm{e}}=\tilde{c}_{\mathrm{e}}^{2}$, obtained from the solution to the undamped and unforced Duffing equation, Eq. 30, for $n_{0}=0.59$ (orange dash), compared with the full-quantum result (blue solid). Also shown are results from MQC dynamics under focused sampling (dotted) and under Wigner sampling (red dash-dotted), obtained using $\Omega=50$.
@@
To assess how closely the MQC Rabi oscillation at $n_{0}=0.59$ follows the quantum result, we present in Fig. 3 a comparison of $P_{\mathrm{e}}=\tilde{c}_{\mathrm{e}}^{2}$ obtained from the Duffing equation, Eq. 30, with that from the quantum Rabi model, Eq. 7, as a function of time. As can be seen, both results are in broad agreement, with a matching dominant oscillation component. The mode-softening effect contained in the Duffing equation manifests as a population bias of the MQC Rabi oscillation towards the excited state of the TLS.
@@
Fig. 3 includes MQC simulation results obtained by directly solving Eq. 16 for $n_{0}=0.59$. In these simulations, $\Omega=50g$ is taken in order to ensure that the RWA is applicable, and that the initial phase of the optical field does not affect the dynamics. The latter justifies an initialization of the classical coordinates at $p=0$ and $q>0$, and negates the need for trajectory sampling. The results coming out of this approach are in close agreement with those produced by the Duffing equation, Eq. 30, substantiating the validity of our analysis.
@@
Our findings suggest that for the model at hand, MQC dynamics are best performed with an initialization of the classical coordinates with a well-defined energy lying slightly above the zero-point energy, by setting $n_{0}=0.59$. This contrasts with the approach commonly taken in MQC dynamics wherein classical coordinates are stochastically sampled from a Wigner quasi-probability distribution Wigner (1932), giving rise to a distribution of initial energies that average out to the zero-point energy. To further investigate this contrast, we have also included in Fig. 3 MQC simulation results where such Wigner sampling is conducted. Accordingly, $p$ and $q$ are stochastically drawn from the normal distribution

### v2

@@ I Introduction
Mixed quantum–classical (MQC) dynamics offers a powerful framework for the simulation and interpretation of various phenomena. Throughout the last few decades, MQC dynamics has established itself as a preferred approach for the low-cost simulation of molecular systems, wherein nuclear motion is represented by classical coordinates while a quantum treatment is reserved for the electronic states C. Tully (1998); Nelson et al. (2014); Subotnik et al. (2016); Wang, Akimov, and Prezhdo (2016); Curchod and Martínez (2018); Crespo-Otero and Barbatti (2018). In recent years, it has been similarly employed for the modeling of materials Nie et al. (2014); Smith and Akimov (2019); Jiang et al. (2021); Lively et al. (2024); Krotz and Tempelaar (2024) and condensed-phase correlated phenomena Petrović, Weber, and Freericks (2024).
@@
Although MQC dynamics is commonly found to provide accurate, or at least predictive, results, its realm of applicability remains poorly understood. According to the Bohr correspondence principle Bohr (1920), the classical approximation taken within MQC dynamics is justified so long as the relevant coordinates attain large quantum numbers. However, MQC dynamics has been shown to reach high accuracy even in cases where this criterion is grossly violated Tempelaar and Reichman (2017); Bondarenko and Tempelaar (2023). Such favorable performance may be explained based on the quasiclassical dynamics literature, where it has been shown that classical equations of motion reproduce quantum dynamics for harmonic potentials and in the short-time limit Heller (1981); Herman and Kluk (1984); Kay (2006). Yet, counterintuitive behaviors may arise once quantum and classical equations of motion are combined within a same framework, as famously exemplified by the Schrödinger cat thought experiment Schrödinger (1935). As an alternative, one may therefore resort to decoherence as a criterion for the applicability of MQC dynamics, meaning that the quantum–classical correlations should remain limited.
@@
While most applications of MQC dynamics involve electron–nuclear coupling, light–matter interactions provide a complementary realm of applicability Sakko, Rossi, and Nieminen (2014); Ruggenthaler et al. (2018); Hoffmann et al. (2019a, b, 2020); Hsieh, Krotz, and Tempelaar (2023a, b); Weight, Li, and Zhang (2023). Notably, the canonical Rabi model, involving a two-level system (TLS) interacting with a single optical mode, was originally proposed within a MQC treatment, invoking a classical approximation for the mode Rabi (1937). While originally investigated under high field intensities and with neglect of feedback of the TLS onto the mode, a fully self-consistent treatment has since been explored Graham and Höhnerbach (1984); A. Cives-Esclop and Sánchez-Soto (1999), as well as behaviors in the quantum–classical crossover region Twyeffort Irish and Armour (2022). However, rapid advances in the areas of quantum optics Reiserer and Rempe (2015) and cavity quantum electrodynamics Ebbesen (2016); Ribeiro et al. (2018); Baranov et al. (2018); Hertzog et al. (2019); Hirai, Hutchison, and Uji-i (2020); Dunkelberger et al. (2022); Mandal et al. (2023); Xiang and Xiong (2024) have generated a particular interest in phenomena where field intensities approach the single-photon limit, as governed by the quantum Rabi model. Much remains to be learned about the applicability of MQC dynamics in this limit, and the ability of the quantum Rabi model to expose general guiding principles for MQC modeling.
@@
Here, we apply MQC dynamics to the quantum Rabi model, specifically considering a TLS in its excited state coupled self-consistently to a single resonant optical mode prepared in its ground state. When represented fully quantum-mechanically, this model is known to exhibit Rabi oscillations manifested as harmonic population fluctuations of the TLS Allen and Eberly (2012). As we show analytically within the rotating wave approximation (RWA), upon adopting a classical representation for the optical mode, population oscillations are instead governed by an undamped and unforced Duffing equation Duffing (1918). This equation involves an anharmonic term contributing a “mode softening effect” to the Rabi oscillations. The solutions to this equation are found to follow the quantum results closely upon (approximately) reinforcing zero-point energy when initializing the classical coordinate. These findings inform on the application of MQC dynamics to coherent phenomena involving small quantum numbers.
@@ II Full-quantum representation
Here, $\hat{q}$ and $\hat{p}$ are coordinate-like operators representing the optical mode, |$\ket{\phi_{\mathrm{g}}}$ and |$\ket{\phi_{\mathrm{e}}}$ refer to the ground and excited states of the TLS, $\mu$ is the associated transition dipole, $\lambda$ denotes the coupling strength between the TLS and the mode, and H.c. is short for Hermitian conjugate.
Figure 1: Schematic of the Rabi model. Shown are the ground and excited states of a two-level system (TLS), denoted |$\ket{\phi_{\mathrm{g}}}$ and |$\ket{\phi_{\mathrm{e}}}$, respectively. Also shown are the lowest two quantum levels of an optical mode, superimposed on a classical harmonic oscillator depiction, involving position and momentum coordinates $q$ and $p$, respectively. The classical optical occupancy, $n=z^{*}z$, is indicated. The coupling between the TLS and the optical mode is denoted as $g$.
The quantum Rabi Hamiltonian, Eq. 1, is commonly formulated by expressing the TLS in terms of Pauli spin matrices, which allows mixed states to be accounted for. Since, however, within MQC dynamics we are exclusively dealing with pure states (possibly accounting for mixed states through an ensemble average over pure states), we instead adopt a ground- and excited-state notation, which compared to the Pauli matrix formalism eliminates one variable, thereby simplifying the forthcoming analysis. Moreover, this notation enforces a connection to quantum optics and quantum electrodynamics where the TLS commonly represents the lowest two states of an atom or other quantum emitter. In both research areas, optical phenomena usually remain within the single-quantum limit, while being amenable to the RWA.
Under these conditions, Eq. 1 can be simplified by first expanding the position-like operator $\hat{q}$ into the optical ladder operators,
@@
and by introducing the effective coupling constant $g\equiv\mu\lambda\sqrt{\Omega/2}$ (see Fig. 1). Application of the RWA, which relies on $\Omega\gg g$, then yields a single-quantum Jaynes–Cummings (JC) Hamiltonian Jaynes and Cummings (1963); Shore and Knight (1993); Larson and Mavrogordatos (2021) of the form
@@
which exposes harmonic Rabi oscillations at a frequency of $2g$.
Upon a replacement of the optical mode by a classical oscillator, transient quantum populations can no longer be conveniently evaluated in the eigenbasis of the quantum Hamiltonian, and are to be evaluated in the physical basis instead. In order to contrast the resulting approach with that taken for the full-quantum representation, it is worth re-evaluating the latter within the physical basis. Accordingly, we introduce the wavefunction of the TLS, which is expanded as
@@
and whose evolution is governed by the time-dependent Schrödinger equation, $i\ket{\dot{\Psi}}=\hat{H}\ket{\Psi}$. This yields dynamical equations for the wavefunction coefficients given by
@@
When $\bar{\epsilon}\gg g$ (which is equivalent to $\Omega\gg g$), the evolution of the wavefunction coefficients will be dominated by phase oscillations at a frequency $\bar{\epsilon}$, due to the diagonal contributions to Eq. 9. Notably, these phase oscillations do not contribute to the population dynamics, which is instead dominated by a slower dynamical component. It proves therefore convenient to extract the slower component by dividing out the phase oscillations. Accordingly, we take
@@ III Mixed quantum–classical representation
We now proceed by taking the classical approximation for the optical mode, starting from the quantum Rabi model. Accordingly, we replace the coordinate-like operators by their classical canonical coordinates in Eq. 1, i.e., $\hat{p}\rightarrow p$ and $\hat{q}\rightarrow q$. This substitution yields a MQC representation in which the optical mode is modeled by a classical harmonic oscillator while its interaction with the TLS remains self-consistent, as depicted in Fig. 1.111Here, we avoid the commonly used “semiclassical” denomination, in order to differentiate such treatment from that of a quasiclassical approach, for which similar terminology is oftentimes used. In the following, it proves convenient to express the optical mode in terms of a complex-valued classical coordinate defined as Krotz, Provazza, and Tempelaar (2021); Miyazaki, Krotz, and Tempelaar (2024)
@@
Accordingly, we arrive at a MQC Hamiltonian of the form
@@
We will now follow an approach similarly to that presented above for the full-quantum representation in the physical basis. First, the dynamical equations for the wavefunction coefficients take the form
@@
As before, we divide out the phase oscillations from the wavefunction coefficients. However, according to Eq. 17, the frequency of these phase oscillations is now dependent on $n$, which itself evolves in time. Accordingly, we take
@@
where we adopted the RWA as we did for the full-quantum representation. From this, it follows that
@@
So long as $\Omega\gg g$ is satisfied, the total combined energy of the TLS and the optical field is well approximated as $E=N\Omega$, with $N=\left(n+\tilde{c}_{\mathrm{e}}^{*}\tilde{c}_{\mathrm{e}}\right)$ representing the sum of the occupancy of the TLS excited state and that of the optical field. Since $E$ is a conserved quantity, so is $N$. If we initialize the TLS in its excited state, i.e., $\tilde{c}_{\mathrm{e}}=1$, and denote the initial optical occupancy as $n_{0}$, we thus have that $n+\tilde{c}_{\mathrm{e}}^{2}=n_{0}+1$. Moreover, initializing $\tilde{c}_{\mathrm{e}}$ as real-valued ensures this coefficient will remain real-valued throughout the dynamics. Consequently, Eq. 29 simplifies to
@@
Eq. 30 takes the form of an undamped and unforced Duffing equation Duffing (1918), with a harmonic term $-(2+n_{0})g^{2}$ and an anharmonic term $2g^{2}$. The latter marks a difference in behavior of MQC dynamics compared to the full-quantum representation, the latter being characterized by purely-harmonic dynamics (cf. Eq. 13). The anharmonicity arises from the self-consistent interaction between the TLS and the optical field under the classical approximation taken for the latter. Specifically, the classical approximation introduces a parametric dependence of the quantum Hamiltonian on the classical coordinate, while the classical equations of motion receive a feedback contribution proportional to $\tilde{c}_{\mathrm{e}}^{2}$, ultimately yielding the $\tilde{c}_{\mathrm{e}}^{3}$ term in Eq. 30. (We note that the anharmonicity disappears when feedback of the TLS onto the classical coordinates is neglected.)
The Duffing equation is ubiquitous in many branches of physics and engineering, and has long been under scrutiny Kovacic and Brennan (2011); Cvetićanin (2013). The particular form given by Eq. 30 is solved in terms of Jacobi elliptic functions, with specific solutions dependent on the initial conditions taken. Insightful simplified forms of these solutions have remained out of reach, but what is known is that the solutions are periodic Abramowitz and Stegun (1968). As such, we find persistent oscillations to be retained under MQC dynamics, showing a dependency on the initial optical occupancy, $n_{0}$. Within this oscillatory motion, the anharmonic term in the Duffing equation yields a reduction of the “restoring force” with increasing amplitude of $\tilde{c}_{\mathrm{e}}$, corresponding to an effective decrease in mode stiffness. For this reason, the anharmonic term is said to exert a “mode softening” effect on the Rabi oscillations.
@@
Incorporation of optical vacuum fluctuations thus requires an initialization satisfying $n_{0}>0$. There are various possibilities for which this criterion can be satisfied. For now, however, we chose to remain agnostic to these possibilities, and instead proceed to systematically and numerically assess the effect of classical coordinate initialization on the Rabi oscillations governed by Eq. 30. Notably, through the dependence of this equation on the initial optical occupancy, $n_{0}=z_{0}^{*}z_{0}$, these oscillations depend on the norm of the initial coordinate $z_{0}$, but not on its phase. This is a direct consequence of the condition $\Omega\gg g$, implying that phase oscillations are infinitely rapid compared to the coherent energy exchange between the TLS and the optical mode.
Figure 2: (a) Transient excited-state population of the TLS, $P_{\mathrm{e}}=\tilde{c}_{\mathrm{e}}^{2}$, obtained from the solutions of the undamped and unforced Duffing equation, Eq. 30. Shown are results as a function of the initial optical occupancy, $n_{0}$. (b) Fourier transform of $P_{\mathrm{e}}$. Asymptotic frequency of the dominant harmonic given by Eq. 34 (black dash) and quantum Rabi oscillation frequency at $\omega=2g$ (red dash) are indicated.
We will proceed to assess the dynamics produced by the Duffing equation in terms of the excited-state population of the TLS, $P_{\mathrm{e}}=\tilde{c}_{\mathrm{e}}^{2}$. Other observables of interest include the TLS ground-state population, which due to normalization follows from the excited-state population as $\tilde{c}_{\mathrm{g}}^{2}=1-\tilde{c}_{\mathrm{e}}^{2}$, and the optical occupancy, which through energy conservation follows as $n=E/\Omega-\tilde{c}_{\mathrm{e}}^{2}$. As such, these observables and their behaviors can trivially be inferred from the excited-state population data presented.
Shown in Fig. 2 (a) is $P_{\mathrm{e}}$ obtained from the solutions to Eq. 30, as a function of time and of $n_{0}$. As anticipated, clear periodic oscillations are observed. Evidently, the oscillation frequency is proportional to $n_{0}$, suggesting that the lack of dynamics in the limit $n_{0}\rightarrow 0$ can be understood as an oscillation with infinite periodicity. The frequency trend is further elucidated in Fig. 2 (b), showing the Fourier transform of $P_{\mathrm{e}}$ as a function of frequency, $\omega$, and initial optical occupancy, $n_{0}$. For each $n_{0}$, the Fourier transform was performed over a duration of $gt=200$, after multiplication of $P_{\mathrm{e}}$ by an exponential with an $1/e$ time of $gt=20$ to avoid clipping and after subtracting a 0.5 baseline, upon which the signal was normalized to the maximum peak. A dominant oscillatory component is observable alongside several minor components that vanish as $n_{0}$ increases. This is suggestive of harmonic oscillatory motion being recovered for $n_{0}\gg 0$. This limit turns out to be amenable to an approximate analytical treatment, providing deeper insights into Rabi oscillations emerging under MQC dynamics.
@@
This asymptote is indicated in Fig. 2, and the dominant component predicted by Eq. 30 is seen to indeed approach this frequency with increasing $n_{0}$.
@@
The asymptotic frequency given by Eq. 34 forms an interesting contrast with the known behavior arising under a full-quantum treatment of the quantum Rabi model under the JC Hamiltonian, for which we have $\omega=2g\sqrt{n_{0}+1}$ (with $n_{0}$ representing the quantum occupancy of the optical field). Interestingly, the MQC Rabi model with neglect of feedback of the TLS on the classical optical field instead yields $\omega=2g\sqrt{n_{0}}$ Allen and Eberly (2012). As such, our self-consistent MQC approach presents itself as an interesting “intermediate” between those limiting cases.
@@
Another curious aspect of Eq. 34 is that it predicts $n_{0}=1/2$ to be the value for which MQC dynamics reproduces quantum Rabi oscillations at $\omega=2g$. For this occupancy, the classical field assumes an energy corresponding to that of vacuum fluctuations. This suggests that for MQC dynamics to reproduce the full-quantum result, the optical field needs to be initialized roughly at the zero-point energy. It should be noted, however, that Eq. 34 was derived assuming $n_{0}\gg 0$, and loses validity at small $n_{0}$ values. Indeed, as seen in Fig. 2 (b), the dominant component predicted by Eq. 30 begins to undershoot the asymptotic frequency with decreasing occupancy, in order to satisfy $\omega\rightarrow 0$ for $n_{0}\rightarrow 0$. Consequently, MQC dynamics is found to reproduce the quantum Rabi oscillation frequency (also indicated in Fig. 2 (b)) most closely at an occupancy exceeding the zero-point value, that is, at approximately $n_{0}=0.59$.
@@
Figure 3: Excited-state population of the TLS, $P_{\mathrm{e}}=\tilde{c}_{\mathrm{e}}^{2}$, obtained from the solution to the undamped and unforced Duffing equation, Eq. 30, for $n_{0}=0.59$ (orange dash), compared with the full-quantum result (blue solid). Also shown are results from mixed quantum–classical (MQC) dynamics under focused sampling (dotted) and under Wigner sampling (red dash-dotted), obtained using $\Omega=50g$.
@@
To assess how closely MQC dynamics reproduces the full-quantum result at $n_{0}=0.59$, we present in Fig. 3 a comparison of $P_{\mathrm{e}}=\tilde{c}_{\mathrm{e}}^{2}$ obtained through Eqs. 7 and 30, as a function of time. As can be seen, both transients are in broad agreement, with a matching dominant oscillation component. Notably, we observe a population bias of the Rabi oscillation towards the excited state of the TLS, which is a consequence of the anharmonic term in Eq. 30 reducing the restoring force with increasing amplitude of $\tilde{c}_{\mathrm{e}}$.
@@
Fig. 3 also includes MQC dynamics results obtained by directly solving Eq. 16 for $n_{0}=0.59$. We note that these results do not adopt the RWA. By setting $\Omega=50g$, however, we ensure that inaccuracies introduced upon an adoption of the RWA, as done in the Duffing equation and the full-quantum representation, are negligible, while also ensuring that the initial phase of the optical field does not affect the dynamics. The latter justifies an initialization of the classical coordinates at $p=0$ and $q>0$, and negates the need for trajectory sampling. The results coming out of this approach are in close agreement with those produced by Eq. 30 substantiating the validity of our analysis.
@@
Our findings suggest that for the model at hand, MQC dynamics are best performed with an initialization of the classical coordinates with a well-defined energy lying slightly above the zero-point energy, by setting $n_{0}=0.59$. This contrasts with the approach commonly taken in MQC dynamics wherein initial classical coordinates are stochastically sampled from a Wigner quasi-probability distribution to represent the vacuum wavefunction of the field Wigner (1932), giving rise to a distribution of initial energies the average of which matches the zero-point energy. To further investigate this contrast, we have also included in Fig. 3 MQC results for which such Wigner sampling is conducted. Accordingly, $p$ and $q$ are stochastically drawn from the normal distribution